OpenStudy (khalilforthewin):

Little Miss Buffet takes all the money from her piggy bank and puts it into a savings account at her local bank. The bank promises an annual interest rate of 2.5% on the balance, compounded semiannually. How much will she have after one year if her initial deposit was $200? A) $205.00 B) $205.01 C) $205.02 D) $205.03

OpenStudy (anonymous):

What do you think it is?

OpenStudy (khalilforthewin):

D or B but im not sure which one

OpenStudy (anonymous):

Neither

OpenStudy (anonymous):

What's 2.5% of 200?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

let's go deeper into that question

OpenStudy (khalilforthewin):

250

OpenStudy (anonymous):

So a % goes up to 100%

OpenStudy (anonymous):

So 2.5% Is like saying 2.5/100 (fraction

OpenStudy (khalilforthewin):

40

OpenStudy (anonymous):

And 200 is twice as much as 100 (obviously) So to figure out what 2.5% of 200 is, you need to double 2.5

OpenStudy (anonymous):

So what's 2.5 x 2?

OpenStudy (khalilforthewin):

5

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Little Miss Buffet takes all the money from her piggy bank and puts it into a savings account at her local bank. The bank promises an annual interest rate of 2.5% on the balance, compounded semiannually. How much will she have after one year if her initial deposit was $200? A) $205.00 B) $205.01 C) $205.02 D) $205.03 Which one adds that ammount that you just stated to 200?

OpenStudy (retireed):

it is compounded semiannually.

OpenStudy (retireed):

It is either B or D. Was a correct response.

OpenStudy (anonymous):

Even if it's semiannual, the answer stays the same as annual

OpenStudy (retireed):

WRONG

OpenStudy (retireed):

Why would they bother saying semi-annually?

OpenStudy (anonymous):

Oh, sorry I forgot there was a formula for this

OpenStudy (anonymous):

A=p(1+rn)nt

OpenStudy (retireed):

That's better.

OpenStudy (khalilforthewin):

so is it B?

OpenStudy (anonymous):

p is the initial amount which is 200 in our case r is the interest rate which is 0.025 always write in decimal form n is the number of compounding a year which is 2 because every 6 month t is the year which is 1

OpenStudy (retireed):

A=p(1+rn)^nt make sure to indicate the exponent

OpenStudy (anonymous):

A=200(1+0.0252)2

OpenStudy (anonymous):

It's not B

OpenStudy (khalilforthewin):

its D

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

205.03

OpenStudy (anonymous):

All the options in my opinions are wrong. \[C = P.(1 + r / 100)^{nt}\] \[\ so C = 200.(1 + .025)^{2} = 200.(1.025)^2 = 210.125???\]

OpenStudy (anonymous):

which does not match any option? Is option or data correct?

OpenStudy (anonymous):

Yeah it is the answer's D

OpenStudy (anonymous):

You're using the wrong formula for this

OpenStudy (anonymous):

A=p(1+rn)nt p is the initial amount which is 200 in our case r is the interest rate which is 0.025 always write in decimal form n is the number of compounding a year which is 2 because every 6 month t is the year which is 1

OpenStudy (retireed):

Your right the equation was wrong A=p(1+r/n)^nt divided the rate by the number of periods

OpenStudy (anonymous):

Wait, so none of the answer's are correct?

OpenStudy (anonymous):

Yep sorry I forgot the r /n thing.

OpenStudy (retireed):

It is D but your equation was incorrect and I missed the mistake.

OpenStudy (anonymous):

No the answer is correct. It was my mistake. I mistook the formular.

OpenStudy (anonymous):

Ahh, ok

OpenStudy (retireed):

You followed the incorrect equation.

OpenStudy (anonymous):

Yeah, because I was pretty sure that D was the answer

OpenStudy (anonymous):

@retirEEd You are right. Its not multiplication with the periods in rate, it should divide

OpenStudy (retireed):

Sorry for the confusion, but that's how we learn not to make that mistake in the future.

OpenStudy (anonymous):

Yeah :)